OpenStudy (anonymous):

Solve x + 2y = 13 4x + 8y = –9

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

they are parallel

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

to solve these, do i just find the lines and then write what the lines are?

OpenStudy (anonymous):

for unique solution, \[\frac{a _{1}}{a _{2}}\neq \frac{b _{1}}{b _{2}}\] for infinite solutions, \[\frac{a _{1}}{a _{2}}=\frac{b _{1}}{b _{2}}=\frac{c _{1}}{c _{2}}\] for no solution, \[\frac{a _{1}}{a _{2}}=\frac{b _{1}}{b _{2}}\neq\frac{c _{1}}{c _{2}}\]